OpenStudy (anonymous):

How do you solve 8+4m=9m-7

OpenStudy (anonymous):

first get m over to the other side 8+ 4m = 9m - 7 12 m = 9m - 7 12 - 9 = 3m -7 3m -7, then solve using m.

OpenStudy (anonymous):

9m-7=8+4m you then subtract 4m from 9m your equation then looks like this 5m-7=8 then you add 7 to 8 your equation looks like this then 5m=15 then divide by 5 5m/5=15/5 m=3